# Determination of the density in the linear elastic wave equation

### v2

@@ 1. Introduction
For the elastic wave equations, the inverse problem is less understood. Rachele [18] proved that the jets of $\lambda,\mu,\rho$ at the boundary can be fully recovered. For interior determination, the boundary control method has not been successfully applied to the full Lamé equation (1). For a Lamé-type equation having the same principal part that admits a full $P/S$-decoupling, the boundary control was used to recover the wavespeeds [3, 4, 12]. Via reduction to geometric inverse problems, the uniqueness of the two wavespeeds $c_{P},c_{S}$ (or equivalently $\frac{\lambda}{\rho},\frac{\mu}{\rho}$) is known (1) assuming $(\Omega,g_{P/S})$ is simple [19]; or (2) assuming that $\partial\Omega$ is convex with respect to $g_{P/S}$ and $(\Omega,g_{P/S})$ admits a strictly convex function [26] (known as the strictly convex foliation condition, see [30, 16] for more discussions on this condition). The proof is based on recovering the lens relations related to the two wavespeeds and the applying the known rigidity results [10, 24]. With $c_{P}$ and $c_{S}$ already determined, under the above two geometric assumptions, [20, 6] study the recovery of the density $\rho$ using the subprincipal term of the $P$-wave amplitude and the inversion of the longitudinal geodesic ray transform of a second order tensor. But the results need to essentially assume a technical assumption $\lambda\neq 2\mu$. These results can be extended to uniqueness of piecewise smooth parameters [8, 7]. We also refer to [9] for a uniqueness result for the density under different assumptions. For a nonlinear elastic wave equation, the density can be recovered utilizing the nonlinearity [31].
@@
In this work we study the recovery of the density $\rho$ using $S$-wave amplitude with known $c_{P}$ and $c_{S}$. With $c_{P/S}$ already known, we use the notation $\Lambda_{\rho}=\Lambda_{\lambda,\mu,\rho}$ for simplicity. The readers will see that the problem reduces to the inversion of the transverse geodesic ray transform of a second order tensor. The invertibility of this type of ray transform is proved recently in [32, 33] under the strictly convex foliation condition. As pointed out in [24] the foliation condition is a natural generalization of the Herglotz [13] and the Wieckert-Zoeppritz [34] conditions, which are typically assumed for the Earth. The main result of this paper is as follows.
@@
Under the strictly convex foliation conditions for both $g_{P}$ and $g_{S}$, we indeed have the the uniqueness of all parameters $\lambda,\mu,\rho$, with preliminary uniqueness results for $\frac{\lambda}{\rho}$ and $\frac{\mu}{\rho}$ in [26]. To be more precise, we have the following corollary.
Corollary 1.
Assume $\partial\Omega$ is strictly convex with respect to $g_{P}$ and $g_{S}$, and $(\Omega,g_{P/S})$ admits a strictly convex function. If $T$ is greater than the length of all geodesics in $(\Omega,g_{P/S})$, then $\Lambda_{\lambda,\mu,\rho}$ determines $\lambda,\mu,\rho$ in $\overline{\Omega}$ uniquely.
Remark 3.
Under the simplicity condition, the injectivity of the related transverse ray transform is also known when the curvature is close to zero [21], and one can directly have uniqueness of $\rho$ under the same assumption for $(\Omega,g_{S})$.
@@ 2. Geometric optics solutions
(5) $\{(t,x,\varrho,\varrho\nabla_{x}\varphi_{P/S});t=\varphi_{P/S}(x),\varrho\neq 0\}.$
The phase functions $\varphi_{P/S}(x)$ solve the eikonel equations
$c^{2}_{P/S}|\nabla_{x}\varphi_{P/S}|^{2}=1.$
@@
$(\tau,\xi)=\varrho(1,\nabla_{x}\varphi_{P/S}).$
@@
$\mathrm{d}s^{2}=g_{\alpha\beta}\mathrm{d}x^{\alpha}\mathrm{d}x^{\beta}+c^{2}_{\bullet}\mathrm{d}\tau^{2}.$
From now on the Greek indices range over the values $1,2$. The Christoffel symbols of the Euclidean metric are then
$\begin{split}&\Gamma^{3}_{\alpha\beta}=-\frac{1}{2}c^{-2}\frac{\partial g_{\alpha\beta}}{\partial\tau},\quad\Gamma^{\beta}_{\alpha 3}=\frac{1}{2}g^{\beta\gamma}\frac{\partial g_{\alpha\gamma}}{\partial\tau},\quad\Gamma^{3}_{\alpha 3}=c^{-1}\frac{\partial c}{\partial x^{\alpha}},\\ &\Gamma^{\alpha}_{33}=-cg^{\alpha\beta}\frac{\partial c}{\partial x^{\beta}},\quad\Gamma^{3}_{33}=c^{-1}\frac{\partial c}{\partial\tau}.\end{split}$
@@
[…]
@@
For $P$ waves, the wave vector $\nabla\varphi$ and the principal amplitude/polarization $\mathbf{a}_{0}$ are parallel; for $S$-waves, they are orthogonal. We will use the subprincipal polarization $\mathbf{a}_{1}$ to recover the density. The explicit construction of $\varphi$, $\mathbf{a}_{0}$ and $\mathbf{a}_{1}$ will be carried out in the following two sections.
@@ 3. P-waves and related results
In this section, we first construct solutions of the form (4) that represent $P$-waves. We will regain the same results obtained in [20, 6], but the calculation is done in a different manner (under curvilinear coordinates versus Cartesian coordinates). For $P$-waves, we take $\varphi=\varphi_{P}$ such that

### v3

@@ 1. Introduction
For the elastic wave equations, the inverse problem is less understood. Rachele [19] proved that the jets of $\lambda,\mu,\rho$ at the boundary can be fully recovered. For interior determination, the boundary control method has not been successfully applied to the full Lamé equation (1). For a Lamé-type equation having the same principal part that admits a full $P/S$-decoupling, the boundary control was used to recover the wavespeeds [3, 4, 13]. Via reduction to geometric inverse problems, the uniqueness of the two wavespeeds $c_{P},c_{S}$ (or equivalently $\frac{\lambda}{\rho},\frac{\mu}{\rho}$) is known (1) assuming $(\Omega,g_{P/S})$ is simple [20]; or (2) assuming that $(\Omega,g_{P/S})$ satisfies the strictly convex foliation condition [28]. Recall that a compact Riemannian manifold $(\Omega,g)$ with strictly convex boundary $\partial\Omega$ satisfies the strictly convex foliation condition if it admits a smooth strictly convex function (see [32, 17] for more discussions on this condition). We note here that a function $h$ is called strictly convex if each level set $h=c$ is strictly convex when viewed from $h>c$. The proof is based on recovering the lens relations related to the two wavespeeds and then applying the known rigidity results [10, 26]. With $c_{P}$ and $c_{S}$ already determined, under the above two geometric assumptions, [21, 6] study the recovery of the density $\rho$ using the subprincipal term of the $P$-wave amplitude and the inversion of the longitudinal geodesic ray transform of a second order tensor. But the results need to essentially assume a technical assumption $\lambda\neq 2\mu$. These results can be extended to uniqueness of piecewise smooth parameters [8, 7]. We also refer to [9] for a uniqueness result for the density under different assumptions. For a nonlinear elastic wave equation, the density can be recovered utilizing the nonlinearity [33].
@@
In this work we study the recovery of the density $\rho$ using $S$-wave amplitude with known $c_{P}$ and $c_{S}$. With $c_{P/S}$ already known, we use the notation $\Lambda_{\rho}=\Lambda_{\lambda,\mu,\rho}$ for simplicity. The readers will see that the problem reduces to the inversion of the transverse geodesic ray transform of a second order tensor. To be more specific, we consider a Riemannian manifold $(\Omega,g)$ with boundary $\partial\Omega$. The transverse geodesic ray transform of a second order tensor $f$ is
$T_{2}f(\gamma,\eta):=\int_{\gamma}\langle f(\gamma(s)),\eta\otimes\eta\rangle_{g}\mathrm{d}s=\int_{\gamma}g^{ik}g^{j\ell}f_{ij}(\gamma(s))\eta_{k}(s)\eta_{\ell}(s)\mathrm{d}s,$
where $\gamma$ runs over any geodesic in $(\Omega,g)$ connecting two boundary points, and $\eta$ is an arbitrary parallel vector field along $\gamma$ and orthogonal to $\dot{\gamma}$. In contrast, the longitudinal geodesic ray transform, encountered in [21, 6], of $f$ is
$I_{2}f(\gamma):=\int_{\gamma}\langle f(\gamma(s)),\dot{\gamma}\otimes\dot{\gamma}\rangle_{g}\mathrm{d}s.$
The injectivity of $I_{2}$ has been studied quite extensively (cf. [22, 24, 11, 16, 29]). For the transverse ray transform, Sharafutdinov [22] proved the injectivity of $T_{2}$ assuming that $(\Omega,g)$ is simple and the curvature is close to zero. Recently, the author and collaborator [34, 35] proved the injectivity of $T_{2}$ under the strictly convex foliation condition.
Proposition 1.
([35]) Assume $\partial\Omega$ is strictly convex and $(\Omega,g)$ admits a smooth strictly convex function. Assume also that the second order tensor $f$ is symmetric. If
$T_{2}f(\gamma,\eta)=0$
for any geodesic $\gamma$ in $\Omega$ with endpoints on $\partial\Omega$, and $\eta$ a parallel vector field orthogonal to $\gamma$. Then $f=0$.
As pointed out in [26] the foliation condition is a natural generalization of the Herglotz [14] and the Wieckert-Zoeppritz [36] conditions, which are typically assumed for the Earth.
We will prove that from the Dirichlet-to-Neumann map one can recover $T_{2}N$ for some tensor $N$ depending on $\rho$, where the Riemannian metric $g=g_{S}$ is induced by the $S$-wavespeed. This is done by the construction and analysis of geometric optics solutions, and the explicit form of $N$ is given in (11). Under the foliation condition, the invertibility of $T_{2}$ (cf. Proposition 1) guarantees that we can recover the symmetrization of $N$, $\mathrm{sym}(N)$. Finally we will then recover $\rho$ from $\mathrm{sym}(N)$ by solving a fourth order elliptic partial differential equation.
The main result of this paper is as follows.
@@
Under the strictly convex foliation conditions for both $g_{P}$ and $g_{S}$, the uniqueness results for $\frac{\lambda}{\rho}$ and $\frac{\mu}{\rho}$ have been obtained in [28]. Actually, only the singular part of the Schwartz kernel of the Dirichlet-to-Neumann map is needed.
Proposition 2.
([28]) Assume $\partial\Omega$ is strictly convex with respect to $g_{P/S}$, and $(\Omega,g_{P/S})$ admits a strictly convex function. If $T$ is greater than the length of all geodesics in $(\Omega,g_{S})$, then the knowledge of the Schwartz kernel of $\Lambda_{\lambda,\mu,\rho}$, up to a smooth function, determines $c_{P/S}$ in $\overline{\Omega}$ uniquely.
With the above proposition, we have the following corollary of Theorem 1.
Corollary 1.
Assume $\partial\Omega$ is strictly convex with respect to $g_{P}$ and $g_{S}$, and $(\Omega,g_{P/S})$ admits a strictly convex function. If $T$ is greater than the length of all geodesics in $(\Omega,g_{P})$ and $(\Omega,g_{S})$, then $\Lambda_{\lambda,\mu,\rho}$ determines $\lambda,\mu,\rho$ in $\overline{\Omega}$ uniquely.
@@
One can also apply the uniqueness result of $T_{2}$ in [22] to obtain the uniqueness of $\rho$ under the assumption that the Riemannian manifold $(\Omega,g_{S})$ is simple and its curvature is close to zero.
@@
The rest of the paper is organized as follows. In Section 2, we outline the procedure for the construction of geometric optics solutions. In Section 3, we use $P$-wave geometric optics solutions to reproduce the results in [21, 6]. In Section 4, we use $S$-wave geometric optics solutions to obtain the main result. The explicit construction of the geometric optics solutions used in Section 3 and Section 4 will be provided in the appendix.
@@ 2. Geometric optics solutions
(5) $\{(t,x,\varrho,-\varrho\nabla_{x}\varphi_{P/S});t=\varphi_{P/S}(x),\varrho\neq 0\}.$
The phase functions $\varphi_{P/S}(x)$ solve the Eikonal equation
(6) $c^{2}_{P/S}|\nabla_{x}\varphi_{P/S}|^{2}=1.$
@@
$(\tau,\xi)=\varrho(1,-\nabla_{x}\varphi_{P/S}).$
@@
$g_{E}=\mathrm{d}s^{2}=g_{\alpha\beta}\mathrm{d}x^{\alpha}\mathrm{d}x^{\beta}+c^{2}_{\bullet}\mathrm{d}\tau^{2}.$
From now on the Greek indices range over the values $1,2$. Throughout the paper, we treat $g_{E},g_{P},g_{S}$ as covariant second order tensors. Under ray coordinates, the Euclidean metric is expressed as $(g_{ij})$ with $g_{3j}=c_{\bullet}^{2}\delta_{3j}$, $i,j=1,2,3$. The matrix $(g^{ij})$ is the inverse matrix of $(g_{ij})$. Then obviously, under the ray coordinates the metric $g_{\bullet}=c_{\bullet}^{-2}g_{E}$ is expressed as $(c_{\bullet}^{-2}g_{ij})$. The Christoffel symbols of the Euclidean metric are
$\begin{split}&\Gamma^{3}_{\alpha\beta}=-\frac{1}{2}c^{-2}\frac{\partial g_{\alpha\beta}}{\partial\tau},\quad\Gamma^{\beta}_{\alpha 3}=\frac{1}{2}g^{\beta\delta}\frac{\partial g_{\alpha\delta}}{\partial\tau},\quad\Gamma^{3}_{\alpha 3}=c^{-1}\frac{\partial c}{\partial x^{\alpha}},\\ &\Gamma^{\alpha}_{33}=-cg^{\alpha\beta}\frac{\partial c}{\partial x^{\beta}},\quad\Gamma^{3}_{33}=c^{-1}\frac{\partial c}{\partial\tau}.\end{split}$
@@
Here $\langle\cdot,\cdot\rangle$ denotes the Euclidean inner product and $|\cdot|$ denote the Euclidean norm. Note that
@@
For $P$ waves, the wave vector $\nabla\varphi$ and the principal amplitude/polarization $\mathbf{a}_{0}$ are parallel. Actually, if we take $\mathbf{a}_{0}=a\nabla\varphi$ for some scalar function $a$, then the equation (8) becomes
$\rho\mathbf{a}_{0}=(\lambda+2\mu)|\nabla\varphi|^{2}\mathbf{a}_{0},$
which is automatically satisfied by choosing $\varphi=\varphi_{P}$ satisfying $c_{P}^{2}|\nabla\varphi_{P}|^{2}=1$ (cf. (6)). For $S$-waves, we take $\langle\mathbf{a}_{0},\nabla\varphi\rangle=0$. Now the equation (8) becomes
$\rho\mathbf{a}_{0}=\mu|\nabla\varphi|^{2}\mathbf{a}_{0},$
which holds true if $\varphi=\varphi_{S}$ satisfies $c_{S}^{2}|\nabla\varphi_{S}|^{2}=1$ (cf. (6)).
Denote
$\begin{split}(P\mathbf{a})_{i}=\rho\partial^{2}_{t}a_{i}-&\partial_{m}(\lambda a_{k;\ell}g^{k\ell}g_{ij}+\mu a_{i;j}+\mu a_{j;i})g^{jm}\\ &+\Gamma_{im}^{n}(\lambda a_{k;\ell}g^{k\ell}g_{nj}+\mu a_{n;j}+\mu a_{j;n})g^{jm}\\ &+\Gamma_{jm}^{n}(\lambda a_{k;\ell}g^{k\ell}g_{ni}+\mu a_{n;i}+\mu a_{i;n})g^{jm}.\end{split}$
Notice that $P$ is indeed the elastic wave operator in curvilinear coordinates. Denote also
$\begin{split}(Q\mathbf{a})_{i}=2\mathrm{i}\rho\partial_{t}a_{i}&-\mathrm{i}\left(\partial_{i}(\lambda a_{k}\varphi_{;\ell}g^{k\ell})+\lambda a_{k}\varphi_{;\ell}g^{k\ell}\partial_{m}g_{ij}g^{jm}+\partial_{m}(\mu a_{i}\varphi_{;j}+\mu a_{j}\varphi_{;i})g^{jm}\right)\\ &-\mathrm{i}\left(\lambda a_{k;\ell}g^{k\ell}\varphi_{;i}+\mu(a_{i;j}+a_{j;i})\varphi_{;m}g^{jm}\right)\\ &+\mathrm{i}\Gamma^{n}_{im}(\lambda a_{k}\varphi_{;\ell}g^{k\ell}g_{nj}+\mu a_{n}\varphi_{;j}+\mu a_{j}\varphi_{;n})g^{jm}\\ &+\mathrm{i}\Gamma^{n}_{jm}(\lambda a_{k}\varphi_{;\ell}g^{k\ell}g_{ni}+\mu a_{n}\varphi_{;i}+\mu a_{i}\varphi_{;n})g^{jm},\end{split}$
and
$\begin{split}(R\mathbf{a})_{i}=-\rho a_{i}+\lambda a_{k}\varphi_{;\ell}g^{k\ell}\varphi_{;i}+\mu a_{j}\varphi_{;m}g^{jm}\varphi_{;i}+\mu\varphi_{;j}\varphi_{;m}g^{jm}a_{i}).\end{split}$
Setting $\mathbf{a}_{-2}=\mathbf{a}_{-1}=0$, the amplitudes $\mathbf{a}_{j}$, $j=0,1,2,\cdots$, can be constructed by solving the equations
$P\mathbf{a}_{j-2}+Q\mathbf{a}_{j-1}+R\mathbf{a}_{j}=0,\quad j=0,1,2,\cdots.$
which is obtained by collecting terms of order $\mathcal{O}(\varrho^{2-j})$ in (7). We will use the subprincipal $\mathbf{a}_{1}$ to recover the density. So we will explicitly construct $\mathbf{a}_{0}$ and $\mathbf{a}_{1}$ in the following two sections. The equation $R\mathbf{a}_{0}=0$ is equivalent to (8). For the construction of $\mathbf{a}_{0}$ and $\mathbf{a}_{1}$, we also need to use the following equations
(9) $(R\mathbf{a}_{1}+Q\mathbf{a}_{0})_{i}=0,$
(10) $(R\mathbf{a}_{2}+Q\mathbf{a}_{1}+P\mathbf{a}_{0})_{i}=0.$
@@ 3. P-waves and related results
In this section, we first construct solutions of the form (4) that represent $P$-waves. We will regain the same results obtained in [21, 6], but the calculation is done in a different manner (under curvilinear coordinates versus Cartesian coordinates). This section is not a part of the proof of the main result of this paper.
For $P$-waves, we take $\varphi=\varphi_{P}$ such that
@@
with $\nabla\varphi(x_{0})=\xi_{0}$, $|\xi_{0}|_{g_{P}}=1$. We can first calculate the principal term of the amplitude. Denote $J^{2}=\det(g_{\alpha\beta})$. Assume $\gamma=\gamma_{x_{0},\xi_{0}}$ be the unit-speed geodesic in $(\Omega,g_{P})$ issued from $x_{0}$ in direction $\xi_{0}$.
Lemma 1.
The principal term of the amplitude can be taken to be
$\mathbf{a}_{0}=A_{P}\frac{\nabla\varphi}{|\nabla\varphi|}$
where $A_{P}$ is a scalar function of the form
$A_{P}=J^{-1/2}c_{P}^{1/2}(\lambda+2\mu)^{-1/2}=J^{-1/2}c_{P}^{-1/2}\rho^{-1/2}.$
The proofs of the above lemma and all following lemmas are given in the appendix. Observe that one can not get any information about $\rho$ in the interior of $\Omega$ from the value of $\mathbf{a}_{0}$ at the boundary, which is also pointed out in [21]. To recover the density in the interior, we need to use the subprincipal term of the amplitude $\mathbf{a}_{1}$, as has been done in [21].
If we denote the scalar function $B_{P}=c_{P}^{3/2}\rho^{1/2}J^{1/2}\langle\mathbf{a}_{1},\nabla\varphi\rangle$, we have the following lemma.
Lemma 2.
$B_{P}=B_{P}(\gamma(s))$ satisfies the following ordinary differential equation along $\gamma$ in $(\Omega,g_{P})$
$2\mathrm{i}\frac{\mathrm{d}B_{P}}{\mathrm{d}s}=M_{ij}\dot{\gamma}^{i}\dot{\gamma}^{j}+C,$
where $M$ is a second order (covariant) tensor of the form
and $C$ does not depend on $\rho$.
The above lemma has already been proved in [21]. In the appendix we give a different proof by carrying out all the calculations in the ray coordinates.
Recall that $g_{P}=c_{P}^{-2}g_{E}$ is the metric induced by the $P$-wavespeed. If the geodesic $\gamma$ connects two points $a\in\partial\Omega$ to $b\in\partial\Omega$, then
$B_{P}(b)-B_{P}(a)=\frac{1}{2\mathrm{i}}\int_{\gamma}(M_{ij}(\gamma(s))\dot{\gamma}^{i}(s)\dot{\gamma}^{j}(s)+C)\mathrm{d}s.$
By carefully analyzing the reflection of waves (with possible mode conversions) (see, for example, [30, 8]) at the boundary or using the argument in [21], one can show that the Dirichlet-to-Neumann map $\Lambda_{\rho}$ determines $B_{P}$ on $\partial\Omega$. Similar case is analyzed for acoustic waves in [31]. Therefore we can recover the longitudinal geodesic ray transform of $M$.
$I_{2}M(\gamma)=\int_{\gamma}M_{ij}(\gamma(s))\dot{\gamma}^{i}(s)\dot{\gamma}^{j}(s)\mathrm{d}s.$
We remark here that the same result was obtained in [21, 6], and then the uniqueness of $\rho$ can be proved assuming, for example, $\lambda=2\mu$ only at isolated points, as well as some geometric conditions. Here one needs to use the injectivity, up to natural obstructions, of the longitudinal geodesic ray transform (cf., for example, [22, 16, 29]) under certain geometric conditions. Indeed, it is proved in [21] that if $I_{2}M_{\rho_{1}}=I_{2}M_{\rho_{2}}$, then
$\begin{split}&(c_{P}^{2}-4c_{S}^{2})\Delta^{2}(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})\\ &-4\frac{c_{P}^{4}-5c_{P}^{2}c_{S}^{2}+8c_{S}^{4}}{c_{P}^{2}-c_{S}^{2}}\Delta\left[\nabla(\log\sqrt{\rho_{1}}+\log\sqrt{\rho_{2}})\cdot\nabla(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})\right]=0.\end{split}$
If $\lambda=2\mu$ only at isolated points, the coefficient $(c_{P}^{2}-4c_{S}^{2})$ of the leading order term in the above equation is degenerate only at those isolated points. Then one can use the boundary determination results $\partial_{\nu}^{k}\rho_{1}=\partial_{\nu}^{k}\rho_{2}$, $k=0,1,2,3$ on $\partial\Omega$ (cf. [19]), and the Unique Continuation Principle (cf. [18]) of elliptic equations to derive $\rho_{1}=\rho_{2}$ in $\overline{\Omega}$.
4. S-waves and proof of the main theorem
In this section, we will use solutions of the form (4) that represent $S$-waves to derive the main result of this paper. We work under the ray coordinates associated with the metric $g_{S}=c_{S}^{-2}\mathrm{d}s^{2}$. Now we take
$|\nabla\varphi|^{2}=c_{S}^{-2}$
with $\nabla\varphi(x_{0})=\xi_{0}$, $|\xi_{0}|_{g_{S}}=1$ and $\langle\mathbf{a}_{0},\nabla\varphi\rangle=0$ (or equivalently $a_{03}=0$). Then the identity (8) is satisfied. Assume $\gamma=\gamma_{x_{0},\xi_{0}}$ be the unit-speed geodesic in $(\Omega,g_{S})$ issued from $x_{0}$ in direction $\xi_{0}$.
Similar to the $P$-waves, we have
Lemma 3.
The principal term of the amplitude can be taken to be
$\mathbf{a}_{0}(\gamma(s))=A_{S}c_{S}\eta(s)$
where $A_{S}$ is a scalar function of the form
$A_{S}=J^{-1/2}c_{S}^{-1/2}\rho^{-1/2}$
and $\eta$ is a parallel (cotangent) vector field along the geodesic $\gamma$ in $(\Omega,g_{S})$ and $\langle\eta,\dot{\gamma}\rangle_{g_{S}}=0$.
Still one can not get any information about $\rho$ in the interior of $\Omega$ from the leading order term of the $S$-wave amplitude at the boundary, as for $P$-waves. We will extract information about $\rho$ in the interior from the subprincipal term $\mathbf{a}_{1}$.
Denote the scalar function
$B_{S}=J^{1/2}c_{S}^{-1/2}\rho^{1/2}\langle\mathbf{a}_{1},\eta\rangle_{g_{S}}=J^{1/2}c_{S}^{3/2}\rho^{1/2}\langle\mathbf{a}_{1},\eta\rangle.$
Here $\mathbf{a}_{1}$ and $\eta$ can both be treated as cotangent vectors. Through tedious calculation we find:
Lemma 4.
$B_{S}=B_{S}(\gamma(s))$ satisfies the following ordinary differential equation along $\gamma$
$2\mathrm{i}\frac{\mathrm{d}B_{S}}{\mathrm{d}s}=\langle N,\eta\otimes\eta\rangle_{g_{S}}+C,$
where $N$ is a second (covariant) order tensor of the form
(11) $\begin{split}N=&c_{S}^{2}\Delta\log\sqrt{\rho}g_{S}+2\nabla^{2}\log\sqrt{\rho}+c_{S}^{2}|\nabla\log\sqrt{\rho}|^{2}g_{S}+4\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\nabla\log\sqrt{\rho}\otimes\nabla\log\sqrt{\rho}\\ &+(\nabla\log\sqrt{\rho}\cdot\nabla c_{S}^{2})g_{S}+\nabla\log\sqrt{\rho}\otimes\left(2\frac{c_{P}^{2}-c_{S}^{2}}{c_{S}^{4}}\nabla c_{S}^{2}+c_{S}^{-2}\nabla c_{P}^{2}-\frac{12}{c_{P}^{2}-c_{S}^{2}}\nabla c_{S}^{2}\right),\end{split}$
and $C$ does not depend on $\rho$.
Here
$\langle N,\eta\otimes\eta\rangle_{g_{S}}=g_{S}^{ik}g_{S}^{j\ell}N_{ij}\eta_{k}\eta_{\ell}=c_{S}^{4}g^{ik}g^{j\ell}N_{ij}\eta_{k}\eta_{\ell}.$
If the geodesic $\gamma$ connects two points $a\in\partial\Omega$ to $b\in\partial\Omega$, then
$B_{S}(b)-B_{S}(a)=\frac{1}{2\mathrm{i}}\int_{\gamma}(N(\gamma(s)),\eta\otimes\eta\rangle_{g_{S}}+C)\mathrm{d}s.$
Similar to the discussion for the $P$-waves, we can recover the transverse ray transform of NN
$T_{2}N(\gamma,\eta)=\int_{\gamma}\langle N(\gamma(s)),\eta\otimes\eta\rangle_{g_{S}}\mathrm{d}s$
from the Dirichlet-to-Neumann map $\Lambda_{\rho}$. Here $\gamma$ is an arbitrary geodesic connecting two boundary points, and $\eta$ could be any parallel vector field along $\gamma$ and orthogonal to $\gamma$ with respect the metric $g_{S}$ (also with respect to the Euclidean metric $g_{E}$). We note here that clearly the transverse ray transform of $N$ is equal to that of the symmetrization of $N$.
Next we switch from ray coordinates to Cartesian coordinates. Under the strictly convex foliation condition, by the invertibility of the transverse ray transform (cf. Proposition 1), we can recover the symmetrization of $N$,
recalling $c_{S}^{2}g_{S}=g_{E}$. Here $I$ is the identity matrix (the Euclidean metric under Cartesian coordinates) and $\otimes^{s}$ denotes the symmetrized product
$u\otimes^{s}v=\frac{1}{2}(u\otimes v+v\otimes u).$
In conclusion, if $\Lambda_{\rho_{1}}=\Lambda_{\rho_{2}}$, we have $\mathrm{Sym}(N_{1})=\mathrm{Sym}(N_{2})$, which can be written as a second order differential equation for $\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}}$,
(12) $\begin{split}&\Delta(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})I+2\nabla^{2}(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})+(|\nabla\log\sqrt{\rho_{1}}|^{2}-|\nabla\log\sqrt{\rho_{2}}|^{2})I\\ &+4\frac{c_{P}^{2}-2c_{S}^{2}}{c_{P}^{2}-c_{S}^{2}}\left(\nabla\log\sqrt{\rho_{1}}\otimes\nabla\log\sqrt{\rho_{1}}-\nabla\log\sqrt{\rho_{2}}\otimes\nabla\log\sqrt{\rho_{2}}\right)\\ &+c_{S}^{-2}(\nabla(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})\cdot\nabla c_{S}^{2})I\\ &+\left(2\frac{c_{P}^{2}-c_{S}^{2}}{c_{S}^{4}}\nabla c_{S}^{2}+c_{S}^{-2}\nabla c_{P}^{2}-\frac{12}{c_{P}^{2}-c_{S}^{2}}\nabla c_{S}^{2}\right)\otimes^{s}(\nabla\log\sqrt{\rho_{1}}-\nabla\log\sqrt{\rho_{2}})\\ =&0\quad\text{in }\Omega.\end{split}$
Applying divergence twice to the above equation, we obtain a fourth order linear differential equation for $\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}}$ of the form
$3\Delta^{2}(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})+\sum_{|\alpha|\leq 3}F_{\alpha}(\rho_{1},\rho_{2})D^{\alpha}(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})=0,\quad\text{in }\Omega,$
by noticing that
$\begin{split}\nabla\cdot\nabla\cdot\left[\Delta(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})I+2\nabla^{2}(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}})\right]=3\Delta^{2}(\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}}),\end{split}$
and the remaining terms in (12) only involves first order derivatives of $\log\sqrt{\rho_{1}}-\log\sqrt{\rho_{2}}$. By the Unique Continuation Principle (cf. [18]) and the boundary determination results $\partial_{\nu}^{k}\rho_{1}=\partial_{\nu}^{k}\rho_{2}$, $k=0,1,2,3$ on $\partial\Omega$ (cf. [19]), we can conclude that
$\rho_{1}=\rho_{2}\quad\text{in }\overline{\Omega}.$
This completes the proof of the main theorem.
Appendix A Proof of lemmas
In this appendix, we provide all the calculations needed for the proofs of Lemmas 1-4.
Proof of Lemma 1.